12. Payments of $100 per quarter are made from June 7, Z through inclusive. If the nominal rate of interest convertible quarterly is 6%. a) Find the present value on September 7,Z-1. b) Find the current value on March 7,Z+ 8. c) Find the accumulated value on June 7, Z 12.

show to work with formulas plz. do not use excel. also plz try to explain clearly as you go. I am very confused about this problem
Answer #1

Solution:

a). Present value on september 7, Z-1 :

=  PV(1.5%,47,-100,,1)/(1.015^3)

= 3256.88

b). Current value on March 7,Z+8 :

=  FV(1.5%,47,-100,,1)(1.015^-16)

= 5403.15

c). Accumulated value on June 7, Z+12 :

= FV(1.5%,47,-100,,1)1.015

= 6959.37